OpenStudy (mysangels):

The local gym offers two different yearly plans: Plan One: $ 115 per year Plan Two: $25 per year plus $5 per visit. A) T = Total cost of gym membership & v = # of gym visits. Write linear equations for both plans. B) Graph Plan 1 and Plan 2 in a coordinate plane. C) After how many visits will the cost of both plans be equal?

OpenStudy (throneofbooks):

A: f(x)=115x+0 and f(x)= 25x+5

OpenStudy (mysangels):

What about v, the number of visits?

OpenStudy (throneofbooks):

Give me a sec :)

OpenStudy (throneofbooks):

C= 18 visits

OpenStudy (mysangels):

Would T = 115 & T = 5v + 25 work?

OpenStudy (mysangels):

I got 18 visits with these equations.

OpenStudy (throneofbooks):

T= 115x+0v (or T=115x) and T=25x+5v T= total amount of money X= number of years V= number of visits.

OpenStudy (jtvatsim):

If you assume that you are only looking at one year, then, yes mysangels, your equations would be fine. If you need to consider many years, then Throneofbooks has a good way of capturing that situation.

OpenStudy (mysangels):

It doesn't specify, so that's what I thought. How would you graph Throneofbooks's equations?

OpenStudy (jtvatsim):

Technically, throneofbooks's equations would need to be plotted in 3D. : ) So, for your assignment, it's probably safe to assume they meant only to think for only one year.

OpenStudy (mysangels):

When I substituted the equations to equal each other, I got 18. But how would you get 18 if the y intercept is 25?

OpenStudy (throneofbooks):

25 multiplied by the number of visits= y intercept

OpenStudy (mysangels):

They equal the y values, not the y intercept

OpenStudy (mysangels):

Ok. I think I get it now. Thank you, guys!